OpenStudy (anonymous):

Solve for the indicated variable: P = 3r + 2s; for s

OpenStudy (e.cociuba):

P=3r + 2s P-3r=2s (P-3r)/2 = s Step 1: Subtract "3r" from both sides of the equation. Step 2: To get "s" by itself, we need to divide both sides by 2. Our final answer is (P-3r)/2 = s

OpenStudy (anonymous):

s= 1/2P - 3r/2 :)